OpenStudy (anonymous):

Last Problem in Proving Identities: Thanks. 2tanx/(1-tan^2x)+1/(2cos^2x-1) = cosx+sinx/(cosx-sinx) ? How?

OpenStudy (anonymous):

\[\frac{ 2tanx }{(1-\tan^2x)} + \frac{ 1 }{ 2\cos^2x-1 }\]

OpenStudy (anonymous):

this is ur LHS??

OpenStudy (anonymous):

LHS? what's that? sorry don't know LHS

OpenStudy (anonymous):

I MEAN THE LEFT HAND SIDE EQUATION OF THAT EQUAL TO SIGN

OpenStudy (anonymous):

y-yes sir/ma'am.